OpenStudy (anonymous):

How to factor 9a^2 + 6ab + b^2 + 6a + 2b - 3?

OpenStudy (mertsj):

\[9a ^{2}+6ab +b ^{2}= (3a+b)^{2} and 6a + 2b = 2(3a+b) \]

OpenStudy (mertsj):

So now you need to replace 3a+b with x and get \[x ^{2}+2x -3\]

OpenStudy (mertsj):

And its factors are (x+3)(x-1) but x = 3a+b so the factors are: (3a+b+3)(3a+b-1)

OpenStudy (anonymous):

Thank you!!